OpenStudy (anonymous):

Input the equation for the following line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. A line with y-intercept of 2 that is parallel to 2x + y = -5.

OpenStudy (texaschic101):

first we need to find the slope of 2x + y = -5 because a parallel line will have the same slope. So we will put the equation in y = mx + b form, where m is the slope. 2x + y = -5 --- subtract 2x from both sides y = -2x - 5 the number in the m position (the slope) is -2 so our parallel line will have a slope of -2. Now we use y = mx + b and since our slope(m) is -2 and our y intercept(b) is 2, then our parallel line will be : y = -2x + 2 Now we just need it in standard form Ax + By = C y = -2x + 2 --- add 2x to both sides 2x + y = 2 <=== standard form...your answer

OpenStudy (anonymous):

thanks